OpenStudy (anonymous):

Let g(x)=2x and h(x)=x^2 +4 (g of g)(3)

OpenStudy (amistre64):

what is g(3)?

OpenStudy (lgbasallote):

12....?

OpenStudy (amistre64):

g(g(3)) is 12 yes

OpenStudy (anonymous):

yes 12 is the answer. i don't know the process though.

OpenStudy (lgbasallote):

g(3) is 2(3) = 6 g(g(3))) = g(6) = 2(6) = 12

OpenStudy (amistre64):

(g o g) is another way to write: g(g(x)) so, we evelauate g(x) when x=g(x) 2(g(x)) = 2(2x) = 4x when x=3 ...

OpenStudy (anonymous):

ok i figured it out thanks!

OpenStudy (amistre64):

:)